# Reducing Task Discrepancy of Text Encoders for Zero-Shot Composed Image Retrieval

### v1

@@ Abstract
Composed Image Retrieval (CIR) aims to retrieve a target image based on a reference image and conditioning text, enabling controllable searches. Due to the expensive dataset construction cost for CIR triplets, a zero-shot (ZS) CIR setting has been actively studied to eliminate the need for human-collected triplet datasets. The mainstream of ZS-CIR employs an efficient projection module that projects a CLIP image embedding to the CLIP text token embedding space, while fixing the CLIP encoders. Using the projected image embedding, these methods generate image-text composed features by using the pre-trained text encoder. However, their CLIP image and text encoders suffer from the task discrepancy between the pre-training task (text $\leftrightarrow$ image) and the target CIR task (image + text $\leftrightarrow$ image). Conceptually, we need expensive triplet samples to reduce the discrepancy, but we use cheap text triplets instead and update the text encoder. To that end, we introduce the Reducing Task Discrepancy of text encoders for Composed Image Retrieval (RTD), a plug-and-play training scheme for the text encoder that enhances its capability using a novel target-anchored text contrastive learning. We also propose two additional techniques to improve the proposed learning scheme: a hard negatives-based refined batch sampling strategy and a sophisticated concatenation scheme. Integrating RTD into the state-of-the-art projection-based ZS-CIR methods significantly improves performance across various datasets and backbones, demonstrating its efficiency and generalizability.
@@ 1 Introduction
Composed Image Retrieval (CIR) is an emerging task aimed at retrieving a target image that closely resembles a reference image while reflecting the changes described in a conditioning text [1]. Using a query composed of image and text allows users to conduct more precise and flexible searches by specifying the desired modifications to the image through text. Supervised CIR methods [2, 3, 4, 5] have been introduced to fuse the information from the bi-modal query, using labeled data in the form of triplets $(I_{r},T_{c},I_{t})$, in which $I_{r}$ is a reference image, $T_{c}$ is a conditioning text, and $I_{t}$ is a target image. However, unlike typical web-crawled image-text datasets [6, 7], acquiring sufficient triplets for training needs expensive manual human annotations. Hence, the existing CIR triplet datasets are typically small, limiting the zero-shot ability of supervised approaches trained on such datasets.
@@
To overcome the dependency on small-scale human-verified triplet datasets, a new task, Zero-Shot Composed Image Retrieval (ZS-CIR), has been recently introduced. The first approach solves this novel task by utilizing the power of recent vision-language (VL) generative models. For example, a line of studies [8, 9, 10, 11] uses text-to-image models like IP2P [12] to synthesize large-scale CIR triplets for supervised model training. Another example can be found in [13], which eliminates the need of training by using image captioning models and large-language models (LLM) during inference. While these approaches achieve decent performance, they are impractical due to their high computational and memory requirements for utilizing generative models. The second approach for removing the dependency on the triplet datasets, which has become the mainstream due to its simplicity, employs an integrable projection module on top of the pre-trained, frozen, and shared VL embedding space, such as CLIP [14]. Namely, a projection module $\phi$ to map a CLIP image embedding to the CLIP text embedding token space can be trained by solely using images [15, 16] or texts [17]. Then, as we illustrated in Figure 1, these methods first project the embedding of the query image to a special text token embedding [\$] using $\phi$. Then, it is combined with the conditioning text [$T_{c}$] to form a prompt “a photo of [\$] that [$T_{c}$]”. Finally, the combined prompt is used as a query for the text-to-image retrieval which utilizes the pre-trained VL embedding space.
@@
Figure 1: The task discrepancy of projection-based ZS-CIR methods [15, 16, 17] between the pre-training task (image-text alignment) and the ZS-CIR task (image-text composition).
@@
The core assumption of the second approach, which often is referred to as projection-based ZS-CIR [15, 16, 17], is that the pre-trained text encoder should be robust enough to combine information from both the projected text token embedding and the conditioning text. However, we argue that this can cause significant task discrepancy for the pre-trained text encoder between the image-text alignment pre-training task and the ZS-CIR task. For example, in Figure 1, if we assume there is an ideal caption that accurately describes the target image, then the target image embedding (Fig. 1c) will align well with the embedding (Fig. 1b) of that caption due to the contrastive learning of the text and image encoders of CLIP. In contrast, in the ZS-CIR task, the text encoder instead receives a concatenated caption that combines the projected text token [\$] and the conditioning text. However, the text encoder typically is not trained to encode complex textual modifications—such as addition, negation, comparison, and spatial relationships—to the reference image, which are common in the conditioning text. As a result, there is no guarantee that the latent textual embedding of the concatenated caption (Fig. 1a) closely aligns with that of the target image embedding (Fig. 1c).
@@
To that end, in this paper, we aim to reduce the task discrepancy of the text encoder only with cheap text triplets. The triplets $(T_{r},T_{c},T_{t})$ can be automatically generated without human labor [1, 18] and intensive resources [8, 9, 10, 11], where $T_{r}$ is the reference caption and $T_{t}$ is a target caption. Using these triplets, we devise a target-anchored text contrastive learning, which trains the text encoder to update the embeddings of the concatenated caption $T_{r+c}$ (formed by simple concatenation of reference caption $T_{r}$ and conditioning text $T_{c}$) to align closely with the fixed embedding of the target caption $T_{t}$, which serves as an anchor point obtained from the frozen text encoder. We also propose two techniques to enhance the effectiveness of such language-only supervision further: a batch sampling strategy that incorporates hard negatives in each mini-batch and a refined concatenation scheme for $T_{r}$ and $T_{c}$ to reduce the training-inference task discrepancy. We note our approach can be seamlessly integrated with existing projection-based ZS-CIR methods [15, 16, 17] by replacing their text encoder with our updated text encoder while fixing other modules, e.g., the image encoder and $\phi$. Moreover, due to the benefits of language-only training, as highlighted by [17], our approach is not only efficient in the dataset generation process, but also in the training process.
Our experimental results demonstrate that our proposed method, dubbed as RTD (Reducing Task Discrepancy of text encoders for Composed Image Retrieval), substantially improves the ZS-CIR performance in diverse evaluation datasets (CIRR [1], CIRCO [16], FashionIQ [18], and GeneCIS [19]). Namely, when integrated into the existing projection-based ZS-CIR methods (SEARLE [16], Pic2Word [15], and LinCIR [17]), RTD consistently enhances performance across two different size backbones (ViT-B/32 and ViT-L/14), underscoring the generality of our approach. We even observe that the integrating RTD into LinCIR/SEARLE (ViT-B/32) has a more significant effect than switching to a larger backbone (ViT-B/32 $\rightarrow$ ViT-L/14). Our systematic ablation analyses reveal that such performance enhancement primarily results from reducing the task discrepancy of the text encoder, rather than merely tuning the textual backbone network with additional data. Instead of updating all parameters of the text encoder, we also show that a more efficient approach, which selectively updates only a few layers of the text encoder, can be effective as well. Moreover, we verify that using template-based text triplets (without LLMs) is sufficient to achieve strong competitive ZS-CIR performance, highlighting the low data acquisition cost of our approach.
@@ 2 Related Work
Projection-based ZS-CIR.
Projection-based CIR methods, such as Pic2Word [15], SEARLE [16] and LinCIR [17], are built upon the frozen CLIP [14] model, which includes a visual encoder $\psi_{V}$ and a text encoder $\psi_{T}$. These methods first project the latent image embedding $v=\psi_{V}(I)$, where $I$ is an input image, to the token embedding space using a projection module $\phi$. Assuming the projected token embedding as a special token [\$], these methods predict the composed embedding by encoding “a photo of [\$] that [$T_{c}$]”, where [$T_{c}$] is the given conditioning text. See Figure 1 for the details of the inference stage of projection-based CIR methods. Each projection-based CIR methods employ a different training scheme for $\phi$. We will explain their differences in Section 4.1. Although they show promising ZS-CIR performances, these methods rely on the pre-trained CLIP visual and text encoders. However, in practice, the target CIR task is different from the pre-training task of CLIP. In this paper, we aim to reduce the task discrepancy between the CLIP pretext task and the CIR task using an efficient language-only training scheme.
Previous attempts to reduce the task discrepancy between the CLIP pretext task and CIR.
Combiner [20] additionally updates the text encoder to minimize the gap between the target caption feature and the summation of the reference image feature and the instruction text feature. However, this approach needs a number of expensive CIR triplets $(I_{r},T_{c},I_{t})$ for training. Our approach uses text-only triplets $(T_{r},T_{c},T_{t})$, cheap and automatically generated. As another example, Chen and Lai [21] synthesizes a triplet of an original image, the corresponding caption, and the masked image, where treating the original image as the target image, the caption as the conditioning text, and the masked image as the reference image. This approach, however, still has a gap between conditioning text (e.g., “change the dog to a cat”) and image caption (e.g., “a dog is jumping to catch a frisbee”); furthermore, it needs the full fine-tuning of the CLIP model, resulting in changing the visual embeddings in the retrieval database. On the other hand, RTD directly uses the instruction texts for training and does not change the target visual encoder, which enables the reuse of pre-extracted CLIP visual embeddings. Lastly, CIReVL [13] reduces the task discrepancy by making a descriptive caption of the composed query using a large captioning model and LLM. Although CIReVL shows great performance without any training, this method needs inefficient and expensive inferences of BLIP [22, 23] and GPT [24]. Furthermore, it needs a well-tuned task-specific prompt by a skilled user. RTD is much more efficient than CIReVL and fully automated without direct human intervention.
@@ 3 Main Method
In this section, we describe our main method, RTD: we first present the intuition and details of the generation of text triplets, then present the learning framework using them.
3.1 Generation of text triplets
As shown in Figure 1, the projection-based ZS-CIR methods may suffer from significant task discrepancy between the CLIP pre-training task and the target CIR task. Instead of directly using expensive CIR triplets $(I_{r},T_{c},I_{t})$, we aim to generate and employ text triplets $(T_{r},T_{c},T_{t})$, which can be cheaply and automatically generated by LLM [8, 12] or rule-based templates [8], to resolve the task discrepancy. The LLM-based text triplet generation process uses a caption $T_{r}$ as an input of the fine-tuned LLM, whose output predicts the corresponding conditioning text $T_{c}$ and the target caption $T_{t}$. The template-based text triplet generation extracts keywords based on the pre-defined rule (e.g., noun) from the given caption $T_{r}$ and randomly alters the keyword from the pre-extracted keyword sets to generate a target caption $T_{t}$. Then, the conditioning text $T_{c}$ is automatically generated from the pre-defined templates (e.g., “change [original keyword] to [altered keyword]”). Our experiment shows that the fine-tuned LLM is not necessary for achieving strong performances; template-based triplets are sufficient. Note that unlike CompoDiff [8] and IP2P [12], which utilize the text generation step as a preliminary phase for subsequent text-to-image generation, we only require the text generation step. Detailed explanations and examples of text triplets are in the Appendix A.2.
@@ 3.2 Target-anchored text contrastive learning
Now, we explain our approach to update the text encoder for mitigating the task discrepancy solely with the generated text triplets $(T_{r},T_{c},T_{t})$. We first assume that there exists a pre-trained projection module $\phi$ obtained by the projection-based ZS-CIR methods [15, 16, 17]. Recall that for a given reference image $I_{r}$ and conditioning text $T_{c}$, the final composed feature is generated by passing the text prompt “a photo of $\phi(\psi_{V}(I_{r}))$ that $T_{c}$” to the text encoder $\psi_{T}$, where $\psi_{V}$ is the visual encoder and $\phi$ is the projection module (See Figure 1). We aim to update the text encoder $\psi_{T}$ to reduce the discrepancy between the pretext task and ZS-CIR task using the text triplets while maintaining $\psi_{V}$ and $\phi$ frozen.
@@
[Target-anchored text contrastive loss] We apply contrastive learning using a paired caption $(T_{r+c},T_{t})$, where $T_{r+c}$ denotes a concatenated caption of reference caption $T_{r}$ and relative caption $T_{c}$. Namely, we let the representation of the concatenated caption closely approximate that of the target caption. However, solely updating the text encoder while fixing the image encoder can break the alignment between image and text encoders. To prevent the issue, we extract the text embedding of $T_{t}$ using the frozen text encoder $\psi_{T}$, while the concatenated caption $T_{r+c}$ is extracted from the learnable text encoder $\psi_{T}^{tr}$, initialized from $\psi_{T}$. Here, we assume that as the target caption $T_{t}$ is a standard caption, a text embedding $\psi_{T}(T_{t})$, is well-aligned with the frozen image embedding space. Following the assumption, we fix the target textual embedding to serve as an anchor point. This approach helps maintain the pre-trained alignment while learning new relationships. As shown in Section 4.4, anchoring the target textual embedding is essential for fine-tuning the text encoder with our objective.
@@
Now, we define our target-anchored text contrastive loss $\mathcal{L}_{TCL}$ using two text encoders: a frozen pre-trained text encoder $\psi_{T}$ and a learnable text encoder $\psi_{T}^{tr}$ which is initialized with $\psi_{T}$. Textual latent embeddings $\tilde{t}_{r+c}$ and $t_{t}$ are extracted from $\psi_{T}^{tr}$ and $\psi_{T}$, respectively. Namely, $\tilde{t}_{r+c}=\psi_{T}^{tr}(E_{w}^{tr}(T_{r+c}))$ and $t_{t}=\psi_{T}(E_{w}(T_{t}))$, where $E_{w}$ is a word embedding layer. We wish to tune $\psi_{T}^{tr}$ to minimize the distance between the concatenated textual embedding $\tilde{t}_{r+c}$ and the target textual embedding $t_{t}$ while maximizing the distance from other textual embeddings within the batch. Therefore, we employ a symmetric InfoNCE loss [25, 26], which is defined as follows:
$\mathcal{L}_{TCL}=$ $\frac{1}{B}\sum^{B}_{k=1}-\log{\frac{e^{(c(\tilde{t}_{r+c}^{k},t_{t}^{k})/\tau)}}{\sum\limits^{B}_{j=1}{e^{(c(\tilde{t}_{r+c}^{k},t_{t}^{j})/\tau)}}+{\sum\limits_{j\neq k}{e^{(c(t_{t}^{k},t_{t}^{j})/\tau)}}}}}-\log{\frac{e^{(c(t_{t}^{k},\tilde{t}_{r+c}^{k})/\tau)}}{\sum\limits^{B}_{j=1}{e^{(c(t_{t}^{k},\tilde{t}_{r+c}^{j})/\tau)}}+{\sum\limits_{j\neq k}{e^{(c(\tilde{t}_{r+c}^{k},\tilde{t}_{r+c}^{j})/\tau)}}}}}$ (1)
@@
where $c(\cdot,\cdot)$ denotes the cosine similarity, $B$ is the batch size, and $\tau$ is a temperature hyperparameter.
@@
[Refined batch sampling strategy for hard negatives] To further enhance the efficacy of updating the text encoder, we devise a simple yet effective batch sampling strategy that incorporates pairs of $(T_{r+c},T_{t})$ and $(T_{r},T_{r})$ within the same batch. Namely, the concatenated caption $T_{r+c}$ and its corresponding reference caption $T_{r}$ are sampled concurrently. Since all other captions in the batch are considered as negatives in Eq. 1, the concatenated caption $T_{r+c}$ and reference caption $T_{r}$ can serve as hard negatives for one another. Moreover, we believe including $(T_{r},T_{r})$ pairs in the contrastive learning helps the learnable text encoder $\psi_{T}^{tr}$ remain closely aligned with the pre-trained encoder $\psi_{T}$.
@@
[Refined concatenation of reference and conditioning texts] A naive concatenation strategy also can suffer from training-inference task discrepancy because we actually use “a photo of [\$] that [$T_{c}$]” for inference. To tackle this issue, rather than simply concatenating the $T_{r}$ and $T_{c}$, we also use the prompt “a photo of [\$] that [$T_{c}$]” for updating the text encoder, where [\$] is obtained by the reference caption $T_{r}$ with the projection module $\phi$. Instead of obtaining a pseudo-word token with latent image embedding $v$, we utilize a textual latent embedding from the reference caption $T_{r}$, i.e., $\phi(t_{r})$. However, [17] showed that naively replacing the image encoder with the text encoder for the input of $\phi$ will suffer from the modality gap [27], a phenomenon where text and image embeddings have a gap between them. Thus, to reduce the potential negative effect of the modality gap, following LinCIR [17], we inject noise sampled from $\text{Unif}(0,1)\times\mathcal{N}(0,1)$ into the textual token representation before it is processed by $\phi$.

### v2

# An Efficient Post-hoc Framework for Reducing Task Discrepancy of Text Encoders for Composed Image Retrieval

@@ Abstract
Composed Image Retrieval (CIR) aims to retrieve a target image based on a reference image and conditioning text, enabling controllable image searches. The mainstream Zero-Shot (ZS) CIR methods bypass the need for expensive training CIR triplets by projecting image embeddings into the text token embedding space, forming a composed query for retrieval. However, we highlight an inherent limitation in these projection-based CIR: a task discrepancy of text encoders between the original pre-training task of the encoders (text $\leftrightarrow$ image) and the target CIR task (image + text $\leftrightarrow$ image), which potentially negatively impacts CIR performance. To reduce such a discrepancy, a naive solution would be to train both image and text encoders with CIR triplets in a supervised manner. Instead, we introduce Reducing Task Discrepancy of Text Encoders (RTD), an efficient text-only post-hoc framework that complements projection-based CIR methods. We devise a novel target-anchored text contrastive learning designed to enhance the capability of the text encoder for CIR. We also propose two key enhancements: (1) a hard negative-based refined batch sampling strategy and (2) a refined concatenation scheme to further mitigate training-inference discrepancy. Integrating RTD into state-of-the-art projection-based methods achieves performance comparable to, or even surpassing, resource-intensive state-of-the-art synthetic CIR triplet-based approaches only with 23 minutes of additional training on 4 A100 GPUs— up to 100×100$\times$100 × faster in training. Our code will be available upon acceptance.
@@ 1 Introduction
Composed Image Retrieval (CIR) aims at retrieving a target image that closely resembles a reference image while reflecting the changes described in a conditioning text. Using a query composed of image and text allows users to conduct more precise and flexible searches by specifying the desired modifications to the image through text. Supervised CIR methods [2, 12, 21] have been introduced to fuse information from bi-modal query, using labeled data from the target domain in the form of triplets $(I_{r},T_{c},I_{t})$, in which $I_{r}$ is a reference image, $T_{c}$ is a conditioning text, and $I_{t}$ is a target image. However, unlike the typical web-crawled image-text datasets [38], acquiring sufficient triplets for training needs expensive manual human annotations. To overcome the dependency on small-scale and human-verified triplets of the target domain, several works utilize the power of recent generative models. For example, a line of studies [16, 45, 22, 48] uses text-to-image models like IP2P [5] or large-language models (LLM) [43, 6] to synthesize large-scale CIR triplets for training, in place of the target-domain CIR triplets. While these methods achieve strong performance, they are often impractical due to the high computational and memory requirements for utilizing generative models.
@@
Another approach for removing the dependency on the CIR triplets, which often is referred to as projection-based ZS-CIR [36, 4, 17, 41, 42, 26], employs an integrable projection module on top of the pre-trained, frozen, and shared VL embedding space, such as CLIP [34]. Namely, a projection module $\phi$, which maps a CLIP image embedding to the CLIP text token embedding space, can be trained by solely using images [36, 4] or texts [17]. During inference, as illustrated in Fig. 1, these methods first project the embedding of the query image to a text token embedding [\$] using the function $\phi$. This embedding is then combined with the conditioning text [$T_{c}$] to create the prompt “a photo of [\$] that [$T_{c}$]”, which is used as a query for the text-to-image retrieval.
@@
Figure 1: The task discrepancy of projection-based ZS-CIR methods between the pre-training task (image-text alignment) and the ZS-CIR task (image-text composition).
@@
The core assumption of the projection-based CIR is that the pre-trained text encoder should be robust enough to combine information from both the projected text token embedding and the conditioning text. However, we argue that this can cause a significant task discrepancy for the pre-trained text encoder between the original image-text alignment pre-training task (of CLIP) and the CIR task. For example, in Fig. 1, consider an ideal caption that accurately describes the target image. Since the CLIP text and image encoders are learned through contrastive learning, we can expect that the target image embedding (Fig. 1c) will align well with the embedding (Fig. 1b) of the ideal caption. In contrast, in the projection-based CIR, the text encoder receives a concatenated caption that combines the projected token [\$] and the conditioning text, not an ideal caption. However, as noted by Baldrati et al. [3], the text encoder is not typically trained to encode complex textual modifications—such as addition, negation, or replacement—to the reference image, which are common in conditioning texts. As a result, there is no guarantee that the textual embedding of the concatenated caption (Fig. 1a) closely aligns with the target image embedding (Fig. 1c), which will undermine the retrieval performance.
@@
To that end, we devise a complementary post-processing approach for existing projection-based CIR methods that reduces the task discrepancy of the text encoder efficiently. Ideally, both image and text backbones would be updated using expensive CIR triplets $(I_{r},T_{c},I_{t})$; however, we achieve this using cheap text triplets and only update the text encoder. These text triplets $(T_{r},T_{c},T_{t})$ consist of a reference caption ($T_{r}$, e.g., “a dog catching a frisbee”), a conditioning text ($T_{c}$, e.g., “change dog to cat”), and a target caption ($T_{t}$, e.g., “a cat catching a frisbee”), respectively. Given that the reference captions $T_{r}$ are readily available from conventional caption datasets, $T_{c}$ and $T_{t}$ can be automatically generated without any human labor [28, 46] or intensive resources [16, 45, 22, 48]. Using these triplets, we introduce a target-anchored text contrastive learning, in which the text encoder updates the embedding of the concatenated caption $T_{r+c}$ of $T_{r}$ and $T_{c}$ (e.g., “a dog catching a frisbee”+ “change the dog to a cat”) to align closely with the fixed embedding of the target caption $T_{t}$ from the frozen text encoder. We further enhance this text-only training approach with two key components: a batch sampling strategy that ensures hard negatives in each mini-batch and a refined concatenation scheme for $T_{r}$ and $T_{c}$ designed to reduce the training-inference task discrepancy caused by our text-only training framework.
Our experimental results demonstrate that our proposed method, dubbed as RTD (Reducing Task Discrepancy of Text Encoders), consistently and considerably improves the performance in diverse evaluation datasets (CIRR [28], CIRCO [4], FashionIQ [46], COCO object composition [36], and GeneCIS [44]), when integrated with existing projection-based CIR methods (SEARLE [4], Pic2Word [36], LinCIR [17], Context-I2W [42], and FTI4CIR [26]). Consequently, compared to resource-intensive state-of-the-art synthetic CIR triplets-based method like MagicLens [48], RTD achieves comparable results, with less than a 1% gap in CIRR R@1 and 0.1% in FashionIQ R@10 when combined with FTI4CIR (ViT-L/14). When using larger or fine-tuned backbones (ViT-G/14 or FSC-CLIP [31]) with LinCIR, RTD achieves superior results than MagicLens in those metrics.
We note that RTD achieves the above results with significantly higher training efficiency in both data generation and training time. Specifically, our rule-based text triplet generation requires less than 10 minutes to construct 1M text triplets—570$\times$ faster than CIR triplet generation from Compodiff [16]. With 4 NVIDIA A100 GPUs, the additional training time for RTD is just 23 minutes for training the ViT-L/14 backbone, making it approximately 10 to 100$\times$ faster than synthetic CIR triplets-based methods. Even with a significantly larger backbone (ViT-G/14), training takes only 2.5 hours, remaining negligible in comparison. Moreover, we investigate the effectiveness of RTD across various text triplet generation strategies and different backbone sizes (ViT-B/32, ViT-L/14, ViT-G/14) and types (FSC-CLIP [31], CLIP [34]), showing its reproducibility and compatibility.
@@ 2 Related Work
Composed Image Retrieval.
Unlike supervised CIR methods [3, 2, 12, 21], projection-based CIR methods [36, 4, 17, 26, 42, 41, 13] are built upon the frozen CLIP model, where a projection module $\phi$ is trained without CIR triplets. Each method employs a different training scheme for different training schemes for $\phi$ (See Sec. 4.1 for details). Several approaches avoid the need for any training by employing interpolation techniques [19], while others leverage powerful yet resource-intensive models such as LLMs and captioners [20, 47]. While our approach is built upon projection-based CIR methods, its training strategy is closely related to another category of CIR methods [16, 45, 22, 48] that use synthetically generated expensive CIR triplets. However, our method stands out by utilizing text triplets and updating only the text encoder, resulting in more efficient training while maintaining strong performance.
Task discrepancy between the CLIP pre-training task and CIR.
Combiner [3] updates the text encoder to minimize the gap between the target caption feature and the sum of the reference image and conditioning text features. FashionERN [9] addresses reference image dominance by introducing a separate branch to amplify the impact of the conditioning text. However, both Combiner and FashionERN require expensive CIR triplets $(I_{r},T_{c},I_{t})$ for training, whereas our approach uses cheap, automatically generated text-only triplets $(T_{r},T_{c},T_{t})$. As another example, Chen and Lai [7] synthesize a triplet of an original image, its caption, and the masked image, treating them as the target image, conditioning text, and reference image, respectively. This approach, however, still has a gap between conditioning text (e.g., “change dog to cat”) and image caption (e.g., “a dog catching a frisbee”); furthermore, it needs the full fine-tuning of the CLIP model, resulting in changing the visual embeddings in the retrieval database. In contrast, RTD directly uses the conditioning texts for training and does not change the target visual encoder, enabling the reuse of pre-extracted CLIP embeddings. Lastly, CIReVL [20] or LDRE [47] reduce task discrepancy by generating descriptive captions of the composed query using a large captioning model [24] and LLM [6]. While effective without any training, they rely on expensive inference steps and require carefully crafted prompts by skilled users. Whereas, RTD is fully automated, human-free, and more efficient during inference.
@@ 3 Main Method
3.1 Obtaining text triplets
To address the task discrepancy, we collect cheaply and automatically generated text triplets $(T_{r},T_{c},T_{t})$, instead of directly using the expensive CIR triplets $(I_{r},T_{c},I_{t})$. Given reference caption $T_{r}$ from conventional caption datasets, conditioning text $T_{c}$ and target text $T_{t}$ can be generated using two strategies: via large language models (LLMs) [16, 5, 22, 45] or, more efficiently, through rule-based templates [16]. We investigate both strategies and demonstrate that RTD consistently improves performance across them.
For the LLM-based strategy, we use the publicly available text triplets by CompoDiff [16], which are employed in our main experiments if not specified. These triplets are generated by taking a given caption $T_{r}$ as an input of the fine-tuned LLM, whose output predicts the corresponding conditioning text $T_{c}$ and the target caption $T_{t}$. Previous works, such as IP2P [5], CoVR [45], and CASE [22], have also explored generating text triplets using LLMs, differing in LLM model types, input data, and fine-tuning strategies. The original purpose of these text triplet generations is to construct CIR triplets $(I_{r},T_{c},I_{t})$, but all these works also release the corresponding text triplets used for their CIR triplet construction. In addition to these publicly available text triplets, we also implement and evaluate an efficient in-context learning generation strategy using LLaMA3-8B [14] without additional fine-tuning. We conduct experiments with all the aforementioned text triplets in Tab. 6 and observe that RTD consistently delivers significant enhancements.
For cheap, rule-based strategy, we can extract a keyword (e.g., “dog”) from $T_{r}$ (“a dog catching a frisbee”) and replace it with a randomly chosen keyword (e.g., “cat”) [16], forming $T_{c}$. The conditioning text is then generated automatically by using pre-defined templates (e.g., “change [original keyword] to [altered keyword]”). Our experiments show that this simple rule-based variant performs similarly to the LLM-based one.
More detailed explanations and examples of each strategy are provided in Sec. A.2, and a comprehensive comparison of generation costs is provided in App. C.
@@ 3.2 Target-anchored text contrastive learning
Now, we explain our approach to update the text encoder for mitigating the task discrepancy solely with the generated text triplets $(T_{r},T_{c},T_{t})$. We first assume that there exists a pre-trained projection module $\phi$ obtained by the projection-based ZS-CIR methods [36, 4, 17]. Recall that for a given reference image $I_{r}$ and conditioning text $T_{c}$, the final composed feature is generated by passing the text prompt “a photo of $\phi(\psi_{V}(I_{r}))$ that $T_{c}$” to the text encoder $\psi_{T}$, where $\psi_{V}$ is the visual encoder and $\phi$ is the projection module (See Fig. 1). We aim to update the text encoder $\psi_{T}$ to reduce the discrepancy between the pretext task and ZS-CIR task using the text triplets while maintaining $\psi_{V}$ and $\phi$ frozen.
@@
Target-anchored text contrastive loss. We apply contrastive learning using a paired caption $(T_{r+c},T_{t})$, where $T_{r+c}$ denotes a concatenated caption of reference caption $T_{r}$ and conditioning caption $T_{c}$. Namely, we let the representation of the concatenated caption closely approximate that of the target caption. However, solely updating the text encoder while fixing the image encoder can break the alignment between image and text encoders. To prevent the issue, we extract the text embedding of $T_{t}$ using the frozen text encoder $\psi_{T}$, while the concatenated caption $T_{r+c}$ is extracted from the learnable text encoder $\psi_{T}^{tr}$, initialized from $\psi_{T}$. Here, we assume that as the target caption $T_{t}$ is a standard caption, a text embedding $\psi_{T}(T_{t})$, is well-aligned with the frozen image embedding space. Following the assumption, we fix the target textual embedding as an anchor point to maintain the pre-trained alignment while learning new relationships. As shown in Sec. 4.3, this anchoring is essential for fine-tuning the text encoder with our objective.
@@
Now, we introduce our target-anchored text contrastive loss $\mathcal{L}_{TCL}$ using two text encoders: a frozen pre-trained text encoder $\psi_{T}$ and a learnable text encoder $\psi_{T}^{tr}$ which is initialized with $\psi_{T}$. Textual latent embeddings $\tilde{t}_{r+c}$ and $t_{t}$ are extracted from $\psi_{T}^{tr}$ and $\psi_{T}$, respectively. Namely, $\tilde{t}_{r+c}=\psi_{T}^{tr}(E_{w}^{tr}(T_{r+c}))$ and $t_{t}=\psi_{T}(E_{w}(T_{t}))$, where $E_{w}$ is a word embedding layer. We aim to tune $\psi_{T}^{tr}$ to minimize the distance between the concatenated textual embedding $\tilde{t}_{r+c}$ and the target textual embedding $t_{t}$ while maximizing the distance from other textual embeddings within the batch. We employ a symmetric InfoNCE loss [8, 11], as follows:
$\mathcal{L}_{TCL}=$ $\frac{1}{B}\sum^{B}_{k=1}-\log\Bigg{(}{\frac{e^{(c(\tilde{t}_{r+c}^{k},t_{t}^{k})/\tau)}}{\sum\limits^{B}_{j=1}{e^{(c(\tilde{t}_{r+c}^{k},t_{t}^{j})/\tau)}}+{\sum\limits_{j\neq k}{e^{(c(t_{t}^{k},t_{t}^{j})/\tau)}}}}}\Bigg{)}$
$-\log\Bigg{(}{\frac{e^{(c(t_{t}^{k},\tilde{t}_{r+c}^{k})/\tau)}}{\sum\limits^{B}_{j=1}{e^{(c(t_{t}^{k},\tilde{t}_{r+c}^{j})/\tau)}}+{\sum\limits_{j\neq k}{e^{(c(\tilde{t}_{r+c}^{k},\tilde{t}_{r+c}^{j})/\tau)}}}}}\Bigg{)}$ (1)
@@
in which $c(\cdot,\cdot)$ denotes the cosine similarity, $B$ is the batch size, and $\tau$ is a temperature.
@@
Refined batch sampling strategy for hard negatives. To further enhance the efficacy of updating the text encoder, we devise a simple yet effective batch sampling strategy that incorporates pairs of $(T_{r+c},T_{t})$ and $(T_{r},T_{r})$ within the same batch. For example, as presented in Fig. 2, a pair such as ($T_{r+c}$: “A dog catching a frisbee + change dog to cat”, $T_{t}$: “A cat catching a frisbee”) is sampled along with its corresponding reference pair ($T_{r}$: “A dog catching a frisbee”, $T_{r}$: “A dog catching a frisbee”) in the same batch. This setup ensures that the concatenated text $T_{r+c}$ and its corresponding reference text $T_{r}$ implicitly act as hard negatives for each other, as their semantics are much more similar ($T_{r+c}$ is derived from $T_{r}$) than those of other randomly sampled texts in the batch. Explicitly distinguishing the embedding of $T_{r+c}$ from that of $T_{r}$ conceptually aligns well with CIR task, as it encourages the model to better capture modifications (in $T_{c}$). Moreover, we believe including $(T_{r},T_{r})$ pairs in the contrastive learning helps the learnable text encoder $\psi_{T}^{tr}$ remain closely aligned with the pre-trained encoder $\psi_{T}$.
@@
Refined concatenation of reference and conditioning texts. As we use “a photo of [\$] that [$T_{c}$]” for inference, a naive concatenation strategy also can suffer from training-inference task discrepancy. To tackle this issue, instead of simply concatenating the $T_{r}$ and $T_{c}$, we use the prompt “a photo of [\$] that [$T_{c}$]” for updating the text encoder, where [\$] is obtained by the reference caption $T_{r}$ with the projection module $\phi$. Instead of obtaining a pseudo-word token with latent image embedding $v$, we utilize a textual latent embedding from the reference caption $T_{r}$, i.e., $\phi(t_{r})$. However, Gu et al. [17] showed that naively replacing the image encoder with the text encoder for the input of $\phi$ will suffer from the modality gap [25], a phenomenon where text and image embeddings have a gap between them. We tackle this issue by injecting random noise into the textual token representation before it is processed by $\phi$, following Gu et al. [17]. More analyses on variations of noise are in Sec. B.7.